OpenStudy (anonymous):

Find the equation of the tangent line to the graph of the given function at the point with the indicated x-coordinate. f(x) = x + 1/x; x = 3

OpenStudy (anonymous):

slope = 1 -1/x^2=8/9 y=8/9 x +c For x =3, y=10/3 10/3=8/9 * 3 +c c=2/3 so, y=8/9 x + 2/3

OpenStudy (anonymous):

or, 9y =8x+6